OpenStudy (anonymous):

What is the third–degree polynomial function such that f(0) = –24 and whose zeros are 1, 2, and 3? A) f(x) = 4x3 – 24x2 + 44x – 24 B) f(x) = 2x3 – 24x2 + 22x – 24 C) f(x) = 4x3 + 24x2 – 44x – 24 D) f(x) = 2x3 + 24x2 – 22x – 24

OpenStudy (anonymous):

You can check your work by knowing that zeros of solution can be written as (x-r) where r is a zero. so f(x)=(x-1)(x-2)(x-3) where f(0)=-24

OpenStudy (anonymous):

sorry i don't really understand.... can you explain a little more please?

OpenStudy (anonymous):

this is a rule, in order for a number to be a zero of the function, (where y is 0) then it should be written in the form of (x-r) where r is the value of x and the product of all of these roots or solutions should be equal to zero.